# Data augmentation with automated machine learning: approaches and performance comparison with classical data augmentation methods

### v2

@@ 4.2.3 Instance-adaptive dynamic search space
Instead of employing a single augmentation strategy, many recent works such as MetaAugment [131], AdaAug [132], InstaAug [133] and [116] have proposed learning adaptive, instant-dependent data augmentation policies from data. These methods are based on the observation that certain transformations may only produce useful augmentations for specific input types but can be harmful when applied globally. For example, whereas the interpretation of the digit "8" is unaffected by flippng, a vertical flipping of the latin symbol “R” may change it to the Cyrillic letter “Я”, thereby corrupting is semantic interpretation. Similarly, a 180o rotation of the Arabic numeral “6” may change its semantic label to “9” whereas the letter "O" is generally agnostic to such a transformation. Therefore, instead of instead of applying the same set of transformation operations on all input samples, instance-adaptive methods [131, 132, 133, 116] learn input specific transformations according to the dataset and task.
@@
These approaches focus on realizing more fine-grained transformation parameterizations that provide flexible augmentations which can be applied to individual instances instead of using coarse transformations that satisfy all samples of the entire dataset. In this formulation, the instance-level augmentations correspond to specific configurations of transformation hyperparameters. MetaAugment [131] utilizes an auxiliary model within an AutoML framework which re-weights input samples in order to learn the application probabilities of different transformation operations for specific input instances. Thus, by learning the probability of applying transformations for different instances, the sub-model essentially predicts useful augmentations that may be specific to the underlying instance. Based on a similar philosophy, the developers of AdaAug [132] propose an adaptive data augmentation technique that, like [134], learns effective augmentation policies in a category (e.g., [134])- and sample (e.g., [132])-dependent way. Their model re-uses extracted features from the input data to map samples to useful transformations, thus adapting the augmentation operations for each instance. The technique is specifically aimed at enhancing the generalization ability of models trained with this augmentation method. The authors claim that data augmentation policies learned by AdaAug can perform well on new datasets without further fine-tuning. Using a slightly different approach, InstaAug [133] learns instance-adaptive augmentation strategies by mapping instances to distributions of transformation operations instead of single transformations, which are relatively coarse. The authors [133] propose to map input samples to transformation distribution parameters in order to learn instance-specific data augmentations policies from training data. Similar to MetaAugment, InstaAug [133] incorporates a dedicated neural network sub-model, so-called augmentation module, to perform the mapping of input instances to desired transformations. In the training process, the augmentation module samples transformations from the set of transformation distributions and applies them to individual input instances in order to generate augmented samples. The resulting augmented samples are fed into a classification network which is then trained to make predictions and, hence, provides information about the quality of the augmentations.
@@ 4.3 NAS –based search space
While most AutoML-based data augmentation methods typically learn effective data augmentation operations for fixed model architectures, an alternative approach to automated data augmentation is to explore the search space for possible neural network architectures that perform well on the given data. These approaches (e.g., ARDA [44], Ref. [136], T-AutoML [135], Tr-AutoML [16]), instead of learning transformation functions to perform explicit data manipulation, are concerned with automatically generating the network itself. The goal of neural architecture search [57] is to automate the process for creating ML models. In essence, it aims to select the best topology (i.e., network structure, including details of synaptic connections) for a given dataset and task. The approach is to take an input dataset and problem specification, specifically, in this case transformation of input images, and generate a model architecture that solves the given problem better than all other architectures for the given dataset. This involves finding hyperparameters that define the model structure (e.g., filter size and configuration, synaptic connection schemes, pooling and sampling details, model width and depth). The procedure is to iterate through model selection, where in each iteration the algorithm generates a different network structure, sets model parameters and validate its effectiveness based on performance on the target task. The process is completed when the best performing model is found. A large number of studies are devoted to this subject. Some NAS-based automated data augmentation approaches have proposed to simultaneously find the best neural network architectures and effective augmentation policies to apply in order to achieve optimal results. The search space in this scenario consists of possible neural architectures, augmentation primitives and tunable hyperparameters. For instance, Kashima et al. in [32] proposed a NAS-based approach that aims to jointly find best NN architectures and effective augmentation policies that optimizes the performance of these architectures. Their approach combined FasterAA [27] and DARTS [138] to jointly learn good neural architectures and the corresponding optimal data augmentation parameters. In this formulation, DARTS provide the set of possible model architectures while FasterAA implements the basic transformation operations that form the basis for generating data augmentation policies. Yang et al. [135] argue that automating only some components of the ML task while hand-crafting others can result in sub-optimal performance. They introduce a transformer-based AutoML framework that aims to automate the entire deep learning pipeline – model architecture, data augmentation, as well as hyperparameter optimization. Thus, their approach jointly formulates (the problems of) neural architecture search (NAS), data augmentation, model training and evaluation. To achieve this, first, candidate architectures and augmentation operations, as well as the associated hyperparameters are represented as a one-dimentional vector. A predictor is then trained to compare different combinations and configurations neural architectures and augmentations while finding the best hyperparameter values.
@@ 5 Optimization of augmentation policies
The optimization process is an iterative procedure whereby the optimization algorithm is executed repeatedly while comparing the results of each iteration until an optimal or a satisfactory result is achieved. Although gradient-based methods have become the de facto approach for optimizing deep neural networks, the discontinuous nature of augmentation operations and their associated hyperparameters make it difficult to apply these techniques out-of-the-box. Heuristic search strategies are therefore the most viable approach for finding effective augmentation policies. classical optimization methods include random search, particle swarm optimization (PSO), ant colony optimization (ACO), simulated annealing, greedy algorithm, genetic algorithms and particle swarm optimization. The techniques have been developed for solving large-scale optimization problems in computer science. The popularity of these techniques stem from their ability to solve complex and ill-formalized AI problems for which analytical search or gradient-based methods are unsuited or do not provide the required level of performance. While a large diversity heuristic search optimization techniques exist, only a small subset of the methods has been applied for the purpose of automating data augmentation. They include reinforcement learning, Bayesian Optimization (BO) and evolutionary computational algorithms. New search space design strategies have also been proposed that allow much simpler search techniques to produce good performance. We discuss these approaches in this section.
@@ 5.1 Reinforcement learning
Reinforcement learning [141] is one of the first optimization techniques to be used in automated data augmentation [17]. It is a classical optimization technique for solving ill-formed problems and has wide applications in machine learning. Indeed, a very large number of automated data augmentation approaches (e.g., AA [17], FastAA [105], AWS [142], PAA [33] Learn2Augment [143], Ref. [116] and RTS [80]) utilize RL for the optimization task. The reinforcement learning approach is based on the principle of an artificial agent interacting with the environment through specific permissible actions and making observations about the outcomes of different actions through rewards and penalties. In the context of data augmentation, the actions are specified as sequences of basic data transformation operations. Observations are the performance results based on a predefined evaluation criterion. Through repeated actions (i.e., application of the specified augmentation operations), the agent finds optimal augmentation strategies that maximize performance over time. In one of the pioneering works on AutoML-based data augmentation, Cubuk et al. [17] utilize a reinforcement learning technique to simultaneously find good augmentations policies and optimal model hyperparameters that produced state-of-the-art performance on several benchmark datasets (see Figures 16 in Subsection 7.2). The authors [17] utilized a recurrent neural network (RNN) as a controller in the RL formulation to dynamically predict the most effective data augmentation policy for a given dataset in the course of training. Their model, AutoAugment [17], searches for the best transformation operations, their probabilities, as well as the respective intensities with the help of the recurrent neural network (RNN)-based controller.
@@
The approach is exceedingly expensive – even on such a relatively small/simple dataset as CIFAR-10, it take about 5,000 GPU hours to train. Consequently, most subsequent works in this direction aim to achieve comparable performance while reducing the computational overheads. FasterAA [27] and AWS [142], for example, propose to improve computational efficiency by employing weight-sharing mechanism within a reinforcement learning framework and achieved comparable results as AA on image classification tasks but with significantly lower computational budget. Patch AutoAugment (PAA) [33] formulates the data augmentation optimization problem as cooperative multi-agent RL problem in which each aent learns an optimal augmentation policy for a small image patch. A multi-agent reinforcement learning algorithm based on decentralized partially observable Markov decision process [144, 145] is then used to learn a joint optimal augmentation policy over the entire image.
@@
Gowda et al. [143] propose a reinforcement learning approach to automatically learn to select the most useful video frames for augmentations on video datasets without actually executing and evaluating these augmentations. The technique incorporates a so-called Semantic Matching sub-module to exploit the natural association activities (i.e., foregrounds) and backgrounds. Augmentations are created by combining a pair of useful video segments, where one segment’s background is mixed with another’s foreground consisting of actor and objects. The approach achieves a performance gain of up to 4.4% over classical augmentation methods.
@@
Although many of the recent RL-based methods have significantly reduced computational cost of the optimization tasks, relative to other search methods, RL techniques are still more expensive for many practical purposes. Consequently, many different approaches are actively being explored.
@@ 5.2 Bayesian Optimization
Because of the enormous computational requirements of RL methods such as AutoAugment and many follow-up works, recent approaches have focused on reducing computational requirements without incurring significant generalization performance penalty. To this end, a large number of techniques [27, 126, 146] have been devised based on Bayesian optimization [147]. BO methods, as opposed to reinforcement learning approaches discussed above, allow to encode domain knowledge by keeping track of, and using past evaluation results to generate better policies in subsequent searches: For these approaches, the search space is formulated as a probability distribution in which subsequent searches are more focused on areas where the likelihood of the best hyperparameters lie. Thus, the BO approach implements a highly efficiency sampling strategy by selecting only the most promising set of hyperparameters to evaluate based on previous calls to the “evaluator”. A simplified workflow of the process is shown in Figure 12.
@@ 5.3 Evolutionary computation algorithms
As an alternative to the reinforcement learning and Bayesian optimization approaches, many works [148] have proposed evolutionary computational techniques [140] for automating the search for effective data augmentation policies. Algorithms based on such techniques utilize specialized operators inspired by biological evolution processes—specifically, natural selection, crossover and mutation—for searching for, and optimizing the performance of useful augmentations in an iterative process during training (see Figure 13). These techniques use the concept of natural selection to ensure that only high performing augmentation policies are maintained from one iteration to the next. The basic idea is to search from an initial pool of possible augmentations and select good transformation operations, and during each iteration of the search process to continually eliminate non-optimal transformations. The augmentations obtained are combined with other augmentations in a process called crossover. Further, the population-based technique periodically induces random changes in one or more augmentation operations of the current set by replacing some of the parameters with random parameters, thereby mutating the affected augmentations. Consequently, new and better set of augmentations are obtained by the process of crossover and mutation. Thus, in contrast to the other black-box optimization approaches discussed earlier that search over static transformation operations, approaches employing evolutionary computation algorithms allow the underlying augmentation operations to be varied and improved in the course of training. The enhancement is achieved by periodically making random changes to the current best policies, and inter-mixing to produce better characteristics. These operation are known in scientific literature as mutation and crossover, respectively. One of the first works to use evolutionary algorithm method for data augmentation was PBA [34]. Cheng et al. proposed a modified version of PBA, Progressive Population Based Augmentation (PPBA) [148] by gradually reducing the search space in the course of training. This augmentation strategy was applied on multimodal data and showed good results.
@@
When composing a search space for this approach, the main focus is to construct an initial set of useful transformations and a corresponding set of policies which can evolve with time. In this case, the augmentations can be composed using a variety of methods. For example, they can be sample space transformation operations (e.g., [34]) or feature space (e.g., [149]). Cheung et al. [149] and propose an evolutionary computation approach to find good augmentation policies composed in the latent space. Poor performing augmentations in the search space are gradually eliminated in the course of training while good ones are retained and enhance over time.
@@ 5.4 Gradient-based methods
Gradient-based automated data augmentation approaches are designed to approximate the gradients of augmentation hyperparameters rather than to iteratively search a discrete space for good augmentation policies. They mainly aim at achieving higher efficency compared with computational methods like RL and evolutionary algorithms
While gradient-based methods have proven to be so universal and well adapted for training deep neural networks to find optimal parameters in modern machine learning algorithms, they are ill-suited for use with out-of-the-box AutoML-based data augmentation formulations. The application of gradient descent methods in solving automated data augmentation problems is constrained by the inherently non-differentiable nature of the search space. This is as a result of problems associated with discrete nature of transformation hyperparameters (e.g., discrete scale and rotation angle specifications), making it difficult to apply simple and effective gradient-based optimization techniques.
@@
Some new methods such as [107, 111] are based on applying reinforcement learning concepts within a connectionist (i.e., network) framework to estimate the gradients of augmentation parameters. Zhang et al. [107] proposed Adversarial AutoAugment (AdvAA) which computes approximate gradients with the aid of REINFORCE algorithm—a general statistical method based on reinforcement learning—proposed by Williams [78]. A similar approach is also used in Online Hyper-parameter Learning for Auto-Augmentation (OHLAA) [111] to estimate the gradient of the loss function with respect to the augmentation policy hyperparameters.
Different from the above methods, some approaches (e.g., [156, 81]) rely on proxy models where augmentation variables can be represented as continuous and differentiable parameters, and used to indirectly optimize the discrete augmentation hyperparameters (through joint training). DADA [156], for example, proposes a continuous formulation of the discrete search space by introducing a continuous reparameterization of the discrete augmentation policy variables using a differentiable neural network as a surrogate model. The method is based on RELAX, one of several techniques proposed by Grathwohl et al. [150] to extend gradient descent to non-differentiable black-box optimization problems. It eliminates the need for continuous relaxation of the discrete variables by utilizing a surrogate neural network, whose parameters can be used to control the augmentation variables through joint optimization with the augmentation policy hyperparameters. Likewise, Shirobokov et al. [81] utilize differentiable surrogate neural networks that allow the optimization of simulator parameters for the synthesis of artificial data that are representative of real physical processes.

### v3

@@ 4.2.3 Instance-adaptive dynamic search space
Instead of employing a single augmentation strategy, many recent works such as MetaAugment [134], AdaAug [135], InstaAug [136] and [119] have proposed learning adaptive, instant-dependent data augmentation policies from data. These methods are based on the observation that certain transformations may only produce useful augmentations for specific input types but can be harmful when applied globally. For example, whereas the interpretation of the digit "8" is unaffected by flippng, a vertical flipping of the latin symbol “R” may change it to the Cyrillic letter “Я”, thereby corrupting its semantic interpretation. Similarly, a 180o rotation of the Arabic numeral “6” may change its semantic label to “9” whereas the letter "O" is generally agnostic to such a transformation. Therefore, instead of applying the same set of transformation operations on all input samples, instance-adaptive methods [134, 135, 136, 119] learn input specific transformations according to the dataset and task.
@@
These approaches focus on realizing more fine-grained transformation parameterizations that provide flexible augmentations which can be applied to individual instances instead of using coarse transformations that satisfy all samples of the entire dataset. In this formulation, the instance-level augmentations correspond to specific configurations of transformation hyperparameters. MetaAugment [134] utilizes an auxiliary model within an AutoML framework which re-weights input samples in order to learn the application probabilities of different transformation operations for specific input instances. Thus, by learning the probability of applying transformations for different instances, the sub-model essentially predicts useful augmentations that may be specific to the underlying instance. Based on a similar philosophy, the developers of AdaAug [135] propose an adaptive data augmentation technique that, like [137], learns effective augmentation policies in a category (e.g., [137])- and sample (e.g., [135])-dependent way. Their model re-uses extracted features from the input data to map samples to useful transformations, thus adapting the augmentation operations for each instance. The technique is specifically aimed at enhancing the generalization ability of models trained with this augmentation method. The authors claim that data augmentation policies learned by AdaAug can perform well on new datasets without further fine-tuning. Using a slightly different approach, InstaAug [136] learns instance-adaptive augmentation strategies by mapping instances to distributions of transformation operations instead of single transformations, which are relatively coarse. The authors [136] propose to map input samples to transformation distribution parameters in order to learn instance-specific data augmentations policies from training data. Similar to MetaAugment, InstaAug [136] incorporates a dedicated neural network sub-model, so-called augmentation module, to perform the mapping of input instances to desired transformations. In the training process, the augmentation module samples transformations from the set of transformation distributions and applies them to individual input instances in order to generate augmented samples. The resulting augmented samples are fed into a classification network which is then trained to make predictions and, hence, provides information about the quality of the augmentations. A comparison of the main search approaches is presented in Table I.
@@ 4.3 NAS –based search space
While most AutoML-based data augmentation methods typically learn effective data augmentation operations for fixed model architectures, an alternative approach to automated data augmentation is to explore the search space for possible neural network architectures that perform well on the given data. These approaches (e.g., ARDA [47], Ref. [139], T-AutoML [138], Tr-AutoML [16]), instead of learning transformation functions to perform explicit data manipulation, are concerned with automatically generating the network itself. The goal of neural architecture search [60] is to automate the process for creating ML models. In essence, it aims to select the best topology (i.e., network structure, including details of synaptic connections) for a given dataset and task. The approach is to take an input dataset and problem specification, specifically, in this case transformation of input images, and generate a model architecture that solves the given problem better than all other architectures for the given dataset. This involves finding hyperparameters that define the model structure (e.g., filter size and configuration, synaptic connection schemes, pooling and sampling details, model width and depth). The procedure is to iterate through model selection, where in each iteration the algorithm generates a different network structure, sets model parameters and validate its effectiveness based on performance on the target task. The process is completed when the best performing model is found. A large number of studies are devoted to this subject. Some NAS-based automated data augmentation approaches have proposed to simultaneously find the best neural network architectures and effective augmentation policies to apply in order to achieve optimal results. The search space in this scenario consists of possible neural architectures, augmentation primitives and tunable hyperparameters. For instance, Kashima et al. in [35] proposed a NAS-based approach that aims to jointly find best NN architectures and effective augmentation policies that optimizes the performance of these architectures. Their approach combined FasterAA [30] and DARTS [141] to jointly learn good neural architectures and the corresponding optimal data augmentation parameters. In this formulation, DARTS provides the set of possible model architectures while FasterAA implements the basic transformation operations that form the basis for generating data augmentation policies. Yang et al. [138] argue that automating only some components of the ML task while hand-crafting others can result in sub-optimal performance. They introduce a transformer-based AutoML framework that aims to automate the entire deep learning pipeline – model architecture, data augmentation, as well as hyperparameter optimization. Thus, their approach jointly formulates (the problems of) neural architecture search (NAS), data augmentation, model training and evaluation. To achieve this, first, candidate architectures and augmentation operations, as well as the associated hyperparameters are represented as a one-dimentional vector. A predictor is then trained to compare different combinations and configurations neural architectures and augmentations while finding the best hyperparameter values.
@@ 5 Optimization of augmentation policies
The optimization process is an iterative procedure whereby the optimization algorithm is executed repeatedly while comparing the results of each iteration until an optimal or a satisfactory result is achieved. Although gradient-based methods have become the de facto approach for optimizing deep neural networks, the discontinuous nature of augmentation operations and their associated hyperparameters make it difficult to apply these techniques out-of-the-box. Heuristic search strategies are therefore the most viable approach for finding effective augmentation policies. classical optimization methods include random search, particle swarm optimization (PSO), ant colony optimization (ACO), simulated annealing, greedy algorithm, genetic algorithms and particle swarm optimization. The techniques have been developed for solving large-scale optimization problems in computer science. The popularity of these techniques stem from their ability to solve complex and ill-formalized AI problems for which analytical search or gradient-based methods are unsuited or do not provide the required level of performance. While a large diversity of heuristic search optimization techniques exist, only a small subset of the methods has been applied for the purpose of automating data augmentation. They include reinforcement learning, Bayesian Optimization (BO) and evolutionary computational algorithms. New search space design strategies have also been proposed that allow much simpler search techniques to produce good performance. We discuss these approaches in this section.
@@ 5.1 Reinforcement learning
Reinforcement learning [144] is one of the first optimization techniques to be used in automated data augmentation [17]. It is a classical optimization technique for solving ill-formed problems and has wide applications in machine learning. Indeed, a very large number of automated data augmentation approaches (e.g., AA [17], FastAA [108], AWS [145], PAA [36] Learn2Augment [146], Ref. [119] and RTS [83]) utilize RL for the optimization task. The reinforcement learning approach is based on the principle of an artificial agent interacting with the environment through specific permissible actions and making observations about the outcomes of different actions through rewards and penalties. In the context of data augmentation, the actions are specified as sequences of basic data transformation operations. Observations are the performance results based on a predefined evaluation criterion. Through repeated actions (i.e., application of the specified augmentation operations), the agent finds optimal augmentation strategies that maximize performance over time. In one of the pioneering works on AutoML-based data augmentation, Cubuk et al. [17] utilize a reinforcement learning technique to simultaneously find good augmentations policies and optimal model hyperparameters that produced state-of-the-art performance on several benchmark datasets (see Figures 16 in Subsection 7.2). The authors [17] utilized a recurrent neural network (RNN) as a controller in the RL formulation to dynamically predict the most effective data augmentation policy for a given dataset in the course of training. Their model, AutoAugment (AA) [17], searches for the best transformation operations, their probabilities, as well as the respective intensities with the help of the recurrent neural network (RNN)-based controller.
@@
The approach is exceedingly expensive – even on such a relatively small dataset as CIFAR-10, it take about 5,000 GPU hours to train. Consequently, most subsequent works in this direction aim to achieve comparable performance while reducing the computational overheads. FasterAA [30] and AWS [145], for example, propose to improve computational efficiency by employing weight-sharing mechanism within a reinforcement learning framework and achieve comparable results as AA on image classification tasks but with significantly lower computational budget. Patch AutoAugment (PAA) [36] formulates the data augmentation optimization problem as cooperative multi-agent RL problem in which each aent learns an optimal augmentation policy for a small image patch. A multi-agent reinforcement learning algorithm based on decentralized partially observable Markov decision process [147, 148] is then used to learn a joint optimal augmentation policy over the entire image.
@@
Gowda et al. [146] propose a reinforcement learning approach to automatically learn to select the most useful video frames for augmentations on video datasets without actually executing and evaluating these augmentations. The technique incorporates a so-called Semantic Matching sub-module to exploit the natural association of activities (i.e., foregrounds) and backgrounds. Augmentations are created by combining a pair of useful video segments, where one segment’s background is mixed with another’s foreground consisting of actor and objects. The approach achieves a performance gain of up to 4.4% over classical augmentation methods.
@@
Although many of the recent RL-based methods have significantly reduced computational cost of the optimization tasks, relative to other search methods, RL techniques are still more expensive for many practical purposes. Consequently, many alternative approaches are actively being explored.
@@ 5.2 Bayesian Optimization
Because of the enormous computational requirements of RL methods such as AutoAugment and many follow-up works, recent approaches have focused on reducing computational requirements without incurring significant generalization performance penalty. To this end, a large number of techniques [30, 123, 149] have been devised based on Bayesian optimization (BO) [150]. BO methods, as opposed to reinforcement learning approaches discussed in the previous subsection, allow to encode domain knowledge by keeping track of, and using past evaluation results to generate better policies in subsequent searches. For these approaches, the search space is formulated as a probability distribution in which subsequent searches are more focused on areas where the likelihood of the best hyperparameters lie. Thus, the BO approach implements a highly efficiency sampling strategy by selecting only the most promising set of hyperparameters to evaluate based on previous calls to the “evaluator”. A simplified workflow of the process is shown in Figure 12.
@@ 5.3 Evolutionary computation algorithms
As an alternative to the reinforcement learning and Bayesian optimization approaches, many works [151] have proposed evolutionary computational techniques [143] for automating the search for effective data augmentation policies. Algorithms based on such techniques utilize specialized operators inspired by biological evolution processes—specifically, natural selection, crossover and mutation—for searching for, and optimizing the performance of useful augmentations in an iterative process during training (see Figure 13). These techniques use the concept of natural selection to ensure that only high performing augmentation policies are maintained from one iteration to the next. The basic idea is to search from an initial pool of possible augmentations and select good transformation operations, and during each iteration of the search process to continually eliminate non-optimal transformations. The augmentations obtained may be combined with other augmentations in a process called crossover. Further, the population-based technique periodically induces random changes in one or more augmentation operations of the current set by replacing some of the parameters with random parameters, thereby mutating the affected augmentations. Consequently, new and better set of augmentations are obtained by the process of crossover and mutation. Thus, in contrast to the other black-box optimization approaches discussed earlier that search over static transformation operations, approaches employing evolutionary computation algorithms allow the underlying augmentation operations to be varied and improved in the course of training. The enhancement is achieved by periodically making random changes to the current best policies, and inter-mixing to produce better characteristics. These operation are known in scientific literature as mutation and crossover, respectively. One of the first works to use evolutionary algorithm method for data augmentation was PBA [37]. Subsequently, Cheng et al. proposed a modified version of PBA, Progressive Population Based Augmentation (PPBA) [151], by gradually reducing the search space in the course of training. This augmentation strategy was applied on multimodal data and showed good results.
@@
When composing a search space for this approach, the main focus is to construct an initial set of useful transformations and a corresponding set of policies which can evolve with time. In this case, the augmentations can be composed using a variety of methods. For example, they can be sample space transformation operations (e.g., [37]) or feature space (e.g., [152]). Cheung et al. [152] propose an evolutionary computation approach to find good augmentation policies composed in the latent space. Poor performing augmentations in the search space are gradually eliminated in the course of training while good ones are retained and enhance over time.
@@ 5.4 Gradient-based methods
Gradient-based automated data augmentation approaches are designed to approximate the gradients of augmentation hyperparameters rather than to iteratively search a discrete space for good augmentation policies. They are mainly aimed at achieving higher efficency compared with computational methods like RL and evolutionary algorithms
While gradient-based methods have proven to be universal and well adapted for training deep neural networks to find optimal parameters in modern machine learning algorithms, they are ill-suited for out-of-the-box AutoML-based data augmentation formulations. The application of gradient descent methods in solving automated data augmentation problems is constrained by the inherently non-differentiable nature of the search space. This is as a result of problems associated with discrete nature of transformation hyperparameters (e.g., discrete scale and rotation angle specifications), making it difficult to apply simple and effective gradient-based optimization techniques.
@@
Some new methods such as [110, 114] are based on applying reinforcement learning concepts within a connectionist (i.e., network) framework to estimate the gradients of augmentation parameters. Zhang et al. [110] proposed Adversarial AutoAugment (AdvAA) which computes approximate gradients with the aid of REINFORCE algorithm—a general statistical method based on reinforcement learning—proposed by Williams [81]. A similar approach is used in Online Hyper-parameter Learning for Auto-Augmentation (OHLAA) [114] to estimate the gradient of the loss function with respect to the augmentation policy hyperparameters.
Different from the above methods, some approaches (e.g., [159, 84]) rely on proxy models where augmentation variables can be represented as continuous and differentiable parameters, and used to indirectly optimize the discrete augmentation hyperparameters (through joint training). DADA [159], for example, proposes a continuous formulation of the discrete search space by introducing a continuous reparameterization of the discrete augmentation policy variables using a differentiable neural network as a surrogate model. The method is based on RELAX, one of several techniques proposed by Grathwohl et al. [153] to extend gradient descent to non-differentiable black-box optimization problems. It eliminates the need for continuous relaxation of the discrete variables by utilizing a surrogate neural network, whose parameters can be used to control the augmentation variables through joint optimization with the augmentation policy hyperparameters. Likewise, Shirobokov et al. [84] utilize differentiable surrogate neural networks that allow the optimization of simulator parameters for the synthesis of artificial data that are representative of real physical processes. We summarize the main strengths and weaknesses of the various optimization techniques covered in this work in Table II.